OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = (-3/4) x-2?

OpenStudy (anonymous):

so here you essentially have a 2 part problem first you need to know what the slope of your y = (-3/4)x-2 is then you need to write y-y1=m(x-x1) for point (12,4) with the slope of the m being perpendicular to the slope of your equation

OpenStudy (anonymous):

try doing a few steps, if you get stuck i'll clarify some more

OpenStudy (anonymous):

y = (-3/4) x-2 slope = -3/4

OpenStudy (anonymous):

Since the Lines shuld be perpen m1*m2 = -1 -3/4 * m2 = -1 m2 = 4/3

OpenStudy (anonymous):

Using ponit Slope Form : (12,4) y-y1=m(x-x1) y-4 = 4/3 ( x- 12) Nw Just Solve

OpenStudy (anonymous):

thanks to both of you, I wish I could give both of y'all medals. This the only topic in Algebra I just CAN'T seem to understand!